Project

General

Profile

Ops/Comm Scheduling meeting : 23/5/2017

Location : CR21 at 14:00 pm

Phone access : New phonebridge number
UK: 08082380274
US: 631.621.5253
Access code: 1779339517

Agenda

  • Actions
    • Macwaters - Re-make dodgy compressor2 remote cable connector - stands
    • Macwaters - Remove old overheated CSW71 to R9 waiting for John Govans and the van - in progress
    • Ajit - Implement MOM/DC state machine override function for all systems; also need documentation and training for operators - documentation and training in progress
    • Macwaters - Report on status of FC psu repair (stands)
      • The spare is installed and has been run into a short
    • Hall Manager - water system change-over procedure (stands) - collar people on Friday
    • EOverton - Status of miceraid5 disk cleanup (stands) now has a script but needs to run it
    • Progress on Hall list List of Hall activities in the shutdown (Colin) DL were down on masse and made significant progress. Need to confirm if any remaining items
    • CWhyte to request survey of TOF1 - requested, next week as Tony off this week
    • SBoyd to ensure we get updates on detector/data quality and on Target performance at Ops meetings
    • AKurup to check which settings are being used by the active Alarm Handlers
    • SBoyd to conscript extra shift cover for the end of next week : Done. Need BLOCs for the 31/1 and 2 of June (i.e. end of cycle)
  • DC update
  • MOM update
  • Performance plots from TOF calibration run (Durga?) OnRec-9297.pdf
  • Update on the tracker saga
  • Hydrogen system update
  • Run plan
    • Tracker may be ready to operate on Friday
      • Is there still Helium in the tracker?
    • LH2 running will not happen this cycle
    • Extraction of LH2 absorber, re-insertion of the FC and establishment of the magnetic channel: Thursday (25May17)—Saturday (27May17);
    • Magnet commissioning: Sunday (28May17)—Monday (29May17); and
    • Empty-channel data taking (Monday 29May17)—Friday (02Jun17).
    • Last training date on Friday. I've asked shifters to meet the MOM on Friday to assess training plans. The MOM should set a time for this meeting.
  • AOB

Minutes

  • Actions
    • Macwaters - Remove old overheated CSW71 to R9 waiting for John Govans and the van - in progress
    • Macwaters - Report on status of FC psu repair (stands)
      • The spare is installed and has been run into a short
    • Hall Manager - water system change-over procedure (stands) - collar people on Friday
    • EOverton - Status of miceraid5 disk cleanup (stands) now has a script but needs to run it
    • Progress on Hall list List of Hall activities in the shutdown (Colin) DL were down on masse and made significant progress. Need to confirm if any remaining items
    • CWhyte to request survey of TOF1 - requested, due Thursday Removing absorber means will need to re-survey CC magnets after the User Cycle.
    • SBoyd to ensure we get updates on detector/data quality and on Target performance at Ops meetings
    • AKurup to check which settings are being used by the active Alarm Handlers
    • SBoyd to conscript extra shift cover for the end of next week : Done. Need BLOCs for the 31/1 and 2 of June (i.e. end of cycle)
      • BLOCs at end of cycle will be shared between Ray and Henry. Note that both Tanaz and Ray/Henry are on day shift, so if they are called out during the night then MICE will need the stand-by shifter
  • New actions
    • CWhyte to request survey of CC Magnets after cycle
    • FDrielsma to investigate low-efficiency channels in EMR x-z
    • Tracker group: decide what should be done to prevent repeat contamination of Tracker Cryostats
  • DC update
    Work is on Hydrogen and Tracker: see below
  • MOM update
    Ed starts as MOM tomorrow
  • Target performance plots: see S19.pdf
    Looks fine - leave it to carry on
  • Performance plots from TOF calibration run (Durga): see OnRec-9297.pdf
    TOF0 - hope to get calibration in the 3 missing slots. TOF2 - waiting for re-calibration process to finish.
    EMR has some low-efficiency channels in the x-z view right where the beam arrives - should investigate.
  • DS fridge saga: Mark and Mike in mtg. with Joerg at the moment, deciding on threshold for remedial action
  • Update on the tracker cryostat water ingress débâcle
    Have cooled cryostat 2 - "only" lost 20 channels. Have dried out cryostat 2 and are checking performance now. Cryostat 4 is cold, needs testing.
    Cryostat 1 has a leak into the cassette volume at the fibres. Appropriate sealant has been delivered and will be applied tomorrow when Craig is back. Should be cured on Thursday and will then start cooling.
    Possibly a batch of bad (i.e. wet) gas was the cause - need a plan to prevent repeat.
    Need to restart tracker volume helium flow on about Friday.
    We (Ed) still need to do the Tracker RF-timing alignment magic.
  • Hydrogen system update
    Managed to liquefy some Neon, but found some control system issues that caused circulation to stop.
    Should increase intensity of shifter monitoring? LH2 meeting at 15:00.
    Need to warm up on Thursday (to remove absorber) then close PRY. Have Geoff Barber and R12 staff on Thu and Fri, may need shifter assistance on Sat. NB no crane driver on Sat.
  • Run plan
    • LH2 running will not happen this cycle
    • Tracker may be ready to operate on Saturday
    • Extraction of LH2 absorber, re-insertion of the FC and establishment of the magnetic channel: Thursday (25May17)—Saturday (27May17);
      • Need to get FC out of shorts and CC into flip modes
    • Magnet commissioning: Sunday (28May17)—Monday (29May17); and
    • Empty-channel data taking (Monday 29May17)—Friday (02Jun17).
  • AOB
    • DSanders - still would like the wedge
    • HNebrensky: have a leaky circuit on the beamstop and an alarm on Aircon unit 3.
    • HNebrensky: MICE caused a lot of trips at end of March run, and then in May run even though we'd just tweaked the beam-bump. Would like to re-tune beam-bump, need to fit round Hall work. Will contact Bryan.
    • HNebrensky: Alarm Handlers have locked up several times failing to re-draw - need PF/AK to give them a kick but also the restarted ALH has to come up in the appropriate state.
      The MOM has already asked AKurup to come out later in the week.
    • HNebrensky: The DS alarm on the shield pressure wasn't followed up promptly, and then last week we had no response an alarm on the Linde fridge itself, and the flashing LH2 beacon may have been ignored for up to 12 hours. No alarm has registered from the A/C unit. We still need to be paying attention

OnRec-9297.pdf (575 KB) Rajaram, Durga, 23 May 2017 13:10

S19.pdf (75.4 KB) Langlands, Joe, 23 May 2017 13:57